### v1

# On the universality of the halo mass function beyond $\Lambda$CDM cosmology

Abstract
The abundance of dark matter haloes as a function of halo mass is a key diagnostic for constraining the cosmological model. The theoretical framework based on excursion set arguments, when applied to an initial Gaussian random field of density fluctuations, predicts universal behaviour for this quantity, when variables are recast in terms of peak height. The great advantage of this, if true, is that it implies one simply needs to accurately simulate only a single cosmological model to build an emulator for any other cosmology of interest. This tantalising possibility has inspired a number of studies over the years. In practice, the diversity of ways for defining haloes has led to a variety of mixed results concerning this issue. In this work, we utilise a suite of high-resolution cosmological $N$-body simulations, to revisit this question for friends-of-friends haloes. We perform our study in the context of the flat, time evolving dark energy model (hereafter $w$CDM), and with simple modifications of the primordial physics afforded through variations of the scalar power spectral index and its possible running. We construct the universal mass function locus from our fiducial simulation (a $\Lambda$CDM model) and emulate this using a linear interpolating function. We then compare this against the loci that we measure for our array of alternate models. We find mass functions that are consistent with universality to within $\lesssim 5\%$ in the fractional difference, with respect to variations of the 8 cosmological parameters that we have considered (2 variations per parameter) and for redshifts $z<7$.
keywords:
methods: numerical – cosmology: large-scale structure of Universe
pubyear: 2024pagerange: On the universality of the halo mass function beyond $\Lambda$CDM cosmologyC
1 Introduction
One of the keystones of modern cosmology is that the majority of gravitating mass in the Universe, some $\sim 85\%$, is in the form of collisionless cold dark matter (hereafter CDM) with the remainder being in the form of the known standard model particles ‘baryons’ and neutrinos (Planck Collaboration et al., 2014). The true physical nature of the dark matter particle remains a mystery (for a review see Bertone, 2010). Primordial fluctuations in the CDM are thought to be sourced after a period of inflation, whereupon they are well described by a Gaussian random field with a power spectrum that has almost scale-invariant potential fluctuations. The wealth of late-time cosmic structures that we observe today, the sheets, filaments, and clusters – the cosmic web – then forms subsequently through the gravitational instability of these primordial density seeds.
One of the commonly used theoretical frameworks to describe the late-time distribution of matter is the halo model. This simply divides all of the matter into a set of distinct units – haloes (see, Cooray & Sheth, 2002; Asgari et al., 2023, for reviews). If one considers the halo model in its basest form, where all particles are considered to belong to a halo, going down to single particle haloes, then it is simply an equivalence class on the matter. However, to elevate this from a tautology and so make useful predictions one requires a model for the abundance of haloes, their evolution with time, their clustering, and their internal density distributions. With these in hand one can make predictions for an astonishingly wide array of cosmological observables (for examples of applications see Cooray & Sheth, 2002; Asgari et al., 2023). The halo model framework has been of great utility for many of the wide field large-scale structure surveys (especially if one encompasses SHAMs and HAMs etc). For example, recently in the Dark Energy Survey it was extensively used in Friedrich et al. (2021) for building covariance models, and for the KiDS survey by Dvornik et al. (2023) who explored the galaxy matter connection, among others. It will no doubt also play an important role in the exploitation of the upcoming surveys like DESI (DESI Collaboration et al., 2016), Euclid (Laureijs et al., 2011) and LSST (Ivezić et al., 2019) surveys. Thus to fully exploit the data from these missions, there is high demand for theoretical models that can accurately predict the abundance of dark matter haloes over a wide range of masses and cosmic time scales to high precision. And doing so for a wide array of viable scenarios beyond the vanilla flavoured vacuum energy dominated CDM model (here after $\Lambda$CDM). Exploring this is one of the aims of this paper.
An analytic form of the halo mass function (HMF) was first derived by Press & Schechter (1974) by means of applying the spherical collapse model to arbitrary points in the initial linear matter density field. They noted that their model predicted that only half of the mass in the Universe was in collapsed haloes, which they corrected in an ad-hoc fashion by multiplication of their final result by 2! Later on, Bond et al. (1991) provided a robust derivation of their formula, which included the missing factor, using excursion set theory. Interestingly, it was found that the Press-Schechter HMF could be rearranged into a universal form, i.e., being completely independent of cosmology and redshift, through a judicious choice of variables. Although the Press-Schechter function failed to accurately predict the number densities of very large and small mass haloes that were measured in cosmological simulations (Lacey & Cole, 1994; Sheth & Tormen, 1999; Jenkins et al., 2001), the possibility of a universal HMF inspired a number of subsequent studies.
The path to a universal HMF is barred by a number of hurdles, one of which is agreeing what the actual definition of a halo is and its associated mass (see White, 2002, for a discussion) (and for a review of methods and codes see Knebe et al., 2011):
  • FoF Haloes: One of the earliest operational definitions was given through the Friends-of-Friends (FoF) algorithm of Davis et al. (1985). This identifies haloes in numerical simulations as all linked particles with separations less than $b\bar{n}^{-1/3}$, where $\bar{n}$ is the mean particle density and $b$ is the linking parameter typically set to $b=0.2$. This halo finder makes no assumptions about the shape of the halo and is computationally efficient, but can suffer from ‘bridging’ between neighbouring groups. Subsequent studies of FoF haloes show that the mass functions of these objects can exhibit, approximate, universal behaviour (Sheth & Tormen, 1999; Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Crocce et al., 2010; Manera et al., 2010; Courtin et al., 2011; Bhattacharya et al., 2011; Juan et al., 2014).
  • SO haloes: Another common halo finder, that is possibly more physically motivated, is the spherical overdensity (SO) method, which identifies haloes as all matter within a sphere whose average density is some fraction $\Delta$ of the background density (Lacey & Cole, 1994). This was inspired by the spherical collapse model, for which, in the Einstein de-Sitter model, after virialisation the average overdensity is found to be $\sim 180\rho_{\rm crit}$, where $\rho_{\rm crit}$ is the critical density. In many studies the characteristic overdensity is typically fixed at $\Delta=200$ for all cosmological model, this is in contrast to the spherical collapse density which varies with the underlying model (Lacey & Cole, 1993; Eke et al., 1996). Subsequent numerical studies of SO haloes show that the mass functions of these haloes departs from universal behaviour (Tinker et al., 2008; Watson et al., 2013; Despali et al., 2016; McClintock et al., 2019; Ondaro-Mallea et al., 2021).
  • Splashback haloes: More recently, another, perhaps even more physically-motivated, model for halo finding involves using the so called splashback radius (More et al., 2015). The splashback radius is located at the place where the accreted matter reaches its first orbital apocentre after turnaround, which corresponds to a sharp drop in the halo density profile. All matter within this radius is then considered to be part of the halo. A subsequent study, has shown that haloes defined using the splashback-radius method are also potentially good candidates in favour of universality (Diemer, 2020).
In summary, these studies have shown that haloes identified through the FoF method, are a good candidate for exploring the universality of the HMF. They also have the practical advantage over the splashback haloes of having a number of computationally efficient algorithms for measuring them in numerical simulations. In this work, we revisit the universality of the HMF with respect to FoF haloes (with $b=0.2$) in the $\Lambda$CDM framework. We also explore a variety of extensions beyond this framework, including time evolving dark energy, and modifications of the primordial physics including running of the primordial power spectral index. We do this by making use of the suite of high-resolution cosmological $N$-body simulations, the Dämmerung Simulations, carried out by Smith & Angulo (2019). Unlike previous relevant studies, we do not assume any existing functional form of the universal HMF. But instead we do assume that a universal locus of the HMF exists, which in practice we take from our fiducial $\Lambda$CDM suite of runs and compare this with results from our variations in the underlying cosmological model.
This paper is organised as follows: In Section §2 we describe the Dämmerung simulation suite used in this work. In §3 we briefly introduce the theoretical background, describe our estimators for the mass function and their associated uncertainties, and how we test the universality with respect to cosmology and redshift. In §4 we present our results for the set of models and redshifts that we consider in our study, and discuss their implications. Finally, in §6 we conclude our findings and discuss future directions.
Table 1: Cosmological parameters used the Dämmerung simulations. Column 1 denotes the simulation series. The next 8 columns denote the 8 cosmological parameters listed in Eq. (1) that define each simulated model: columns (2) and (3) give $w_{0}$ and $w_{a}$ which are parameters used to define the dark energy equation of state; column (4) gives the density parameter for dark energy $\Omega_{\rm DE}$; columns (5) and (6) give the physical densities of CDM $\omega_{\rm c}=\Omega_{\rm c}h^{2}$ and baryons $\omega_{\rm b}=\Omega_{\rm b}h^{2}$; columns (7), (8) and (9) give the primordial amplitude of scalar perturbations $A_{\rm s}$, the spectral index $n_{\rm s}$ and the running of the spectral index $\alpha_{\rm s}$. Columns (10)-(12) give $\Omega_{\rm m}$, $h$ and $\sigma_{8}$, all of which are derived parameters. Parameter values that are deviations from the fiducial model are highlighted in bold.
Para $w_{0}$ $w_{a}$ $\Omega_{\rm DE}$ $\omega_{\rm c}$ $\omega_{\rm b}$ $A_{\rm s}[\times 10^{-9}]$ $n_{\rm s}$ $\alpha_{\rm s}$ $\Omega_{\rm m}$ $h$ $\sigma_{8}$
Fiducial $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $0.6777$ $0.8284$
V1 $\mathbf{-1.1}$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.8477
V2 $\mathbf{-0.9}$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.8052
V3 $-1.0$ $\mathbf{0.2}$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.8136
V4 $-1.0$ $\mathbf{-0.2}$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.8400
V5 $-1.0$ $0.0$ $\mathbf{0.7429}$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ ${\bf 0.2571}$ $\mathbf{0.7407}$ 0.8438
V6 $-1.0$ $0.0$ $\mathbf{0.6429}$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $0.0$ ${\bf 0.3571}$ $\mathbf{0.6285}$ 0.8137
V7 $-1.0$ $0.0$ $0.6929$ $\mathbf{0.1248}$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $\mathbf{0.6918}$ 0.8580
V8 $-1.0$ $0.0$ $0.6929$ $\mathbf{0.1129}$ $0.02216$ $2.148$ $0.9611$ $0.0$ $0.3071$ $\mathbf{0.6633}$ 0.7981
V9 $-1.0$ $0.0$ $0.6929$ $0.1189$ $\mathbf{0.02327}$ $2.148$ $0.9611$ $0.0$ $0.3071$ $\mathbf{0.6804}$ 0.8216
V10 $-1.0$ $0.0$ $0.6929$ $0.1189$ $\mathbf{0.02105}$ $2.148$ $0.9611$ $0.0$ $0.3071$ $\mathbf{0.6750}$ 0.8352
V11 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $\mathbf{2.363}$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.8688
V12 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $\mathbf{1.933}$ $0.9611$ $0.0$ $0.3071$ $0.6777$ 0.7858
V13 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $\mathbf{1.0092}$ $0.0$ $0.3071$ $0.6777$ 0.8438
V14 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $\mathbf{0.9130}$ $0.0$ $0.3071$ $0.6777$ 0.8136
V15 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $\mathbf{0.01}$ $0.3071$ $0.6777$ 0.8304
V16 $-1.0$ $0.0$ $0.6929$ $0.1189$ $0.02216$ $2.148$ $0.9611$ $\mathbf{-0.01}$ $0.3071$ $0.6777$ 0.8263
2 Numerical simulations
We start with introducing the cosmological model of the Dämmerung simulations, followed by the specifications of the simulation settings. Dämmerung simulations are a suite of dark-matter-only (DMO) cosmological $N$-body simulations carried out by Smith & Angulo (2019), which assume a flat, dark energy dominated cold dark matter model (hereafter $w$CDM), characterised by 8 parameters as follows:
$p_{\alpha}=\{w_{0},w_{a},\Omega_{\rm DE},\omega_{\rm c},\omega_{\rm b},A_{\rm s},n_{\rm s},\alpha_{\rm s}\}\ \ ,$ (1)
where $w_{0}$ and $w_{a}$ define the dark energy equation of state parameter $w(a)$ via
$w(a)=w_{0}+w_{a}(1-a),$ (2)
as proposed by Chevallier & Polarski (2001). $\Omega_{\rm DE}$ is the present day density parameter for dark energy, $\omega_{\rm c}=\Omega_{\rm c}h^{2}$ and $\omega_{\rm b}=\Omega_{\rm b}h^{2}$ are the physical densities of cold dark matter and baryons, and $h$ is the dimensionless Hubble parameter.
The matter power spectrum is initialised by specifying the primordial power spectrum of curvature perturbations and we make use of the following form (Komatsu et al., 2009; Planck Collaboration et al., 2014):
$\Delta^{2}_{\mathcal{R}}(k)=A_{\rm s}\left(\frac{k}{k_{\rm p}}\right)^{(n_{\rm s}-1)+\alpha_{\rm s}\log(k/k_{\rm p})/2}\ ,$ (3)
where $A_{\rm s}$ is the primordial amplitude, $n_{\rm s}$ and $\alpha_{\rm s}$ are the spectral index and the running of the spectral index, all of which are determined at the pivot scale $k_{\rm p}$. The running of the spectral index can also be equivalently written as:
$\alpha_{\rm s}\equiv\frac{{\rm d}n_{\rm s}}{{\rm d}\log{k}}\bigg{|}_{k=k_{\rm p}}.$ (4)
Hence, the matter power spectrum can be written in terms of primordial quantities as:
$P_{\rm m}(k,a)=\frac{8\pi^{2}}{25}\frac{a^{2}g^{2}(a_{\rm early},a)}{\Omega_{\rm m}^{2}}\frac{c^{4}}{H_{0}^{4}}\,T^{2}(k,a)\,k\Delta^{2}_{\mathcal{R}}(k)\ ,$ (5)
where $g(a_{\rm early},a)$ is the growth suppression factor from an early epoch $a_{\rm early}$ to $a$, $T(k,a)$ is the matter transfer function at epoch $a$ and $c/H_{0}$ gives the Hubble scale today. The simulations adopt a fiducial cosmology that is consistent with the Planck Collaboration et al. (2014) cosmological parameters – see Table1 for details.
Note that for this parameterisation, taken along with the assumption of flatness, we can derive a number of other useful cosmological parameters: the matter density parameter $\Omega_{\rm m}=1-\Omega_{\rm DE}$; the dimensionless Hubble rate $h=\sqrt{\left[\omega_{\rm b}+\omega_{\rm c}\right]/\Omega_{\rm m}}$; the square-root of the variance of matter fluctuations in spheres of radius $R=8\ h^{-1}{\rm Mpc}$ can also be computed $\sigma_{8}$ – for details of this see Eq. (15). Through varying each of the 8 cosmological parameters in Eq. (1) positively and negatively, additional 16 variational models are included to investigate the dependence of the nonlinear structure formation on these parameters. Note that these variations were chosen to facilitate the construction of derivatives of the observables with respect to the cosmological parameters and so enable Taylor expansions of the model around the fiducial point to be constructed. The exact values of the cosmological parameters used in all models are listed in Table 1. Note that here we correct typos for the parameters that were presented in Table 1 of Smith & Angulo (2019) for models listed as V5 and V6, which varied the dark energy density (or matter density).
All simulation runs were performed using a special version of the parallel Tree-PM code Gadget-3, which was developed for the purpose of generating the Millennium-XXL simulation (Springel, 2005; Angulo et al., 2012). Each run was performed with $N_{\rm p}=2048^{3}$ dark matter particles and $N_{\rm grid}=2048^{3}$ PM grid in a box of comoving size $L_{\rm box}=500\ h^{-1}\rm{Mpc}$. The particle mass for all simulations was $m_{\rm p}=1.240\times 10^{9}\ h^{-1}M_{\odot}$, with the exception of models V5 and V6 (see Table 1), which have $m_{\rm p}=1.038\times 10^{9}\ h^{-1}M_{\odot}$ and $1.442\times 10^{9}\ h^{-1}M_{\odot}$, respectively, arising due to the variation of $\Omega_{\rm m}$. Note that we here correct also the typos of the particle masses listed in Table 2 of Smith & Angulo (2019). The softening length was set to $l_{\rm soft}=7.5h^{-1}\rm kpc$, which is about $3\%$ of the mean inter-particle spacing $l=L_{\rm box}/\sqrt[3]{N_{\rm p}}$, which is widely used as the empirical optimal softening length (Power et al., 2003).
The fiducial model consists of 10 realisations with the same initial power spectrum but different realisations of the Gaussian random field. These were generated using a modified version of the public 2LPT algorithm of (Crocce et al., 2006). By contrast, only one run was performed for each of the variational models, and the initial Gaussian random field generation was set for each variation to match that of run 1 in the fiducial model with the exception of the power spectrum model and linear growth rates which were modified accordingly for each model. This was done to help minimise the effect of large-scale cosmic variance when making comparisons between variations. Note that we refer to run 1 as the fiducial run hereafter unless otherwise stated. The haloes were identified using the on-the-fly FoF algorithm built into Gadget-3 and was run with the linking length parameter set to $b=0.2$. 63 snapshots were output between $z=49$ and $z=0$ with a hybrid linear-logarithmic time spacing. In this study we primarily focus on 8 snapshots selected at $z=0.00$, $0.51$, $0.99$, $1.50$, $2.05$, $3.08$, $4.54$, and $6.75$.
3 Halo mass functions and universality
3.1 Background theory
In this subsection we define some key theoretical concepts from excursion set theory (for a review see Zentner, 2007). To begin, let us start by defining the number density of haloes with masses between $m$ and $m+{\rm d}m$ at redshift $z$, $n(m,z){\rm d}m$ where $n(m,z)$ is usually called the differential mass function. The mass density of all haloes with masses between $m$ and $m+{\rm d}m$ is then given by $mn(m,z){\rm d}m$. Hence, the fraction of the total mass in the Universe locked up in these dark matter haloes can then be defined as:
$f(\nu){\rm d}\nu\equiv\frac{mn(m,z){\rm d}m}{\overline{\rho}_{0}}\ ,$ (6)
where $\overline{\rho}_{0}$ is the comoving mean matter density of the Universe. Note that we have written the left-hand-side of this equation in terms of the variable $\nu$. In this study we follow the notation in Sheth et al. (2001) and define this to be $\nu=\delta^{2}_{\rm c}(z)/\sigma^{2}(m)$  [^1] One should be careful here on interpreting the literature as some other authors define $\nu$ as $\delta_{\rm c}(z)/\sigma(m)$, which leads to an increase of $\nu f(\nu)$ by a factor of 2 (Note, with a small misuse of terminology, we shall at times refer to this quantity as the ‘peak-height’). Here $\delta_{\rm c}(z)$ is the time evolving, linear theory, critical overdensity for spherical collapse of an object at redshift $z$ and $\sigma^{2}(m)$ is the mass variance of the density field smoothed by the spherical top-hat window function at the present day. Note that for CDM power spectra there is a bijective relation between $\nu$ and $M$ and so one can map back and forth between them interchangeably.
If we were to invoke the ‘halo model’ framework, then we impose the condition that all of the mass in the Universe should be locked up in haloes. This in turn means that if we integrate over all haloes then the fraction of mass in haloes should give unity:
$\int_{0}^{\infty}f(\nu){\rm d}\nu=\int_{0}^{\infty}\frac{mn(m,z){\rm d}m}{\overline{\rho}_{0}}=1\ .$ (7)
We further note that by rearranging Eq. (6), to give the fraction of mass per logarithmic mass interval we can also define the multiplicity function through [^2] Note that this is not really the multiplicity function, as this is actually given by $m^{2}n(m)/\rho_{0}$, which gives the fraction of mass density in haloes per unit range of $\log m$. However we will persevere with this slight mislabelling for convenience. :
$\nu f(\nu)=\frac{m}{\overline{\rho}_{0}}[mn(m,z)]\frac{{\rm d}\log{m}}{{\rm d}\log{\nu}},\ .$ (8)
The first analytic form for the HMF was developed by Press & Schechter (1974), with the derivation being put on a more firm footing by Bond et al. (1991). They found that $n(m,z)$ could be written:
$n_{\rm PS}(m,z)=\sqrt{\frac{2}{\pi}}\frac{\overline{\rho}_{0}}{m^{2}}\frac{\delta_{\rm c}(z)}{\sigma(m)}\bigg{|}\frac{{\rm d}\log{\sigma}}{{\rm d}\log{m}}\bigg{|}\exp{\bigg{[}-\frac{\delta_{\rm c}^{2}(z)}{2\sigma^{2}(m)}\bigg{]}}\ ,$ (9)
which, when recast in terms of the multiplicity function given by Eq. (8) and making use of $\nu$, leads to the simple functional form:
$\nu f_{\rm PS}(\nu)=\sqrt{\frac{\nu}{2\pi}}\exp{\bigg{(}-\frac{\nu}{2}\bigg{)}}\ .$ (10)
This implies a universal behaviour because the variables that are dependent on the cosmology and redshift are inserted into $\nu$. Any change in the cosmology or redshift would only vary $\nu$ rather than $\nu f(\nu)$.
Lacey & Cole (1994) and later Sheth & Tormen (1999) subsequently found that the Press-Schechter model overestimated the abundance of low-mass haloes and underrepresented the abundance of high-mass haloes when compared to numerical simulations. This inspired Sheth & Tormen (1999) and Sheth et al. (2001) to explore ellipsoidal collapse and more complicated excursion set scenarios. This opened the path to more complex possibilities for the analytic form of the HMF. However, when cast in terms of the multiplicity function, they retained the universal structure, e.g. Sheth & Tormen (1999) proposed:
$\nu f(\nu)=A\bigg{[}1+\frac{1}{(a\nu)^{p}}\bigg{]}\sqrt{\frac{a\nu}{2\pi}}\exp{\bigg{(}-\frac{a\nu}{2}\bigg{)}},$ (11)
where $a$ and $p$ were related to the ellipsoidal collapse scenario, but were taken as free parameters, and $A$ can be determined by requiring $\int f(\nu){\rm d}\nu=1$. This function was shown to give a significantly better analytic approximation to simulated data. This approach paved the way for a wide array of numerical studies of the HMF (Jenkins et al., 2001; White, 2002; Reed et al., 2003; Warren et al., 2006; Reed et al., 2007; Lukić et al., 2007; Tinker et al., 2008; Crocce et al., 2010; Bhattacharya et al., 2011; Courtin et al., 2011; Reed et al., 2013; Watson et al., 2013; Castorina et al., 2014; Bocquet et al., 2016; McClintock et al., 2019; Euclid Collaboration et al., 2023)  [^3] Note that we view studies that have chosen to model the mass function using a function $f(\sigma^{-1})$ la Jenkins et al. (2001) as being approximately equivalent to those modelling $\nu$, since $\delta_{\rm c}$ is a very weak function of cosmology and since they encode the same information, but see Courtin et al. (2011) for further discussion of this issue.
As noted earlier, unlike in previous studies, rather than search for a new functional form for the mass function, our aim is to verify whether such behaviour exists and if so whether it holds across a range of cosmological models that includes extensions beyond $\Lambda$CDM.
3.2 Ingredients
In the previous section we defined some important variables and here we specify them in more detail as this will be needed in what follows.
  • For the time evolving, linear theory, critical overdensity for spherical collapse $\delta_{\rm c}$ we take this to be given by the approximation for $\Lambda$CDM as found by (Kitayama & Suto, 1996):
    $\delta_{\rm c}=\frac{3}{20}(12\pi)^{2/3}[1+0.0123\log_{10}\Omega_{\rm m}(z)]\ .$ (12)
    In the above expression the matter density parameter evolution is given by
    $\Omega_{\rm m}(z)=\frac{\Omega_{\rm m,0}(1+z)^{3}}{\Omega_{\rm m,0}(1+z)^{3}+(1-\Omega_{\rm m,0})}\ ,$ (13)
    where $\Omega_{\rm m,0}$ is the matter density parameter at the present time. Note that while we evaluate this expression at the necessary redshift, we do not refer to this as the time dependent collapse threshold. Instead we define this following the excursion set theory approach and scale this linear theory collapse density to the required redshift using the linear theory growth factors. This then gives us the barrier for collapse to be (note that this is done instead of evolving $\sigma(m)$ with time):
    $\delta_{\rm c}(z)=\delta_{\rm c}\frac{D(z=0)}{D(z)}\ ,$ (14)
    where $D(z)$ is the linear growth factor (note that it does not have to be normalised because we only need its ratio here). We have checked that the values of $\delta_{\rm c}$ for the $w$CDM cosmology differ weakly (within $1\%$) from the values given by Eq. (12), regarding the cosmological parameters used in V1, V2, V3 and V4 (see Percival, 2005; Batista, 2021). Thus we apply the same formula to compute $\delta_{\rm c}$ for all models. The computation of $D(z)$ for $w$CDM cosmology is done using the publicly available code CCL (Chisari et al., 2019).
  • We compute the variance of matter fluctuations $\sigma^{2}(m)$ smoothed by a spherical top-hat window function, at the present day using the expression:
    $\sigma^{2}(m)=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}P_{\rm m}(k)W^{2}(kR){\rm d}k,$ (15)
    where $P_{\rm m}(k)$ is the present day linear matter power spectrum given by Eq. (5). The top-hat filter in the Fourier space is
    $W(y)=3\left[\sin{y}-y\cos y\right]/y^{3}\ \ ;\ \ y\equiv kR\ \ ,$ (16)
    where the filter scale $R$ and the mass $m$ enclosed in the sphere are related by $m=4\pi R^{3}\overline{\rho}_{0}/3$. $\overline{\rho}_{0}$ can be simply obtained via the critical density $\rho_{\rm c,0}=2.7754\times 10^{11}h^{-1}M_{\odot}h^{3}{\rm Mpc}^{-3}$ and the matter density parameter $\Omega_{\rm m,0}$ at present. Note that in computing the numerical integral, we set the lower limit to be $k_{\rm min}=2\pi/L_{\rm box}$, since in the simulations there is no power on scales larger than the box wavemode. For the upper limit we take this to be $k_{\rm max}=\pi/l_{\rm soft}$ where $l_{\rm soft}$ is the softening length. We have checked that the integral has converged to much better than sub-percent precision for these limits.
3.3 Estimate of the binned mass function
We now describe how we go about estimating the HMF ${\rm d}n(m)/{\rm d}\log m$ and the multiplicity function $\nu f(\nu)$, as given by Eq. (8), from our simulations.
We start by estimating the mass function and we do this by setting up 8 mass bins per decade in logarithmic scale, i.e., $\Delta\log_{10}m=0.125$. We then compute the mass of each halo by simply multiplying the number of particles in a linked FoF group by the mass per particle: $M_{\rm FoF}=Nm_{\rm p}$. Next we discard the lowest mass haloes from our analysis, due to the following considerations. Warren et al. (2006) showed that the FoF halo mass estimates suffer from a systematic bias, which arises due to the fact that haloes are resolved by discrete particles. They found that if a halo is sampled by too few particles, then the halo masses tend to be biased high on average. Warren et al. (2006) proposed an empirical formula to correct the particle counts, which has the form
$N_{\rm cor}=N(1-N^{-0.6})\ ,$ (17)
where $N_{\rm cor}$ is the estimate of the corrected particle counts and $N$ is the actual number of particles sampling the halo. This formula has been widely used in the literature, especially when treating haloes sampled by several tens of particles (for a recent application to the Millennium TNG simulations see Hernández-Aguayo et al., 2023). Subsequent studies by Lukić et al. (2007) and Courtin et al. (2011) have also found that Eq. (17) is not a general recipe for mitigating this bias, since the correction appears to depend on other factors such as halo concentration, the simulation code used, and the underlying cosmological model. Hence, its application should be checked case by case.
Owing to the fact that to correct for this bias robustly would require a suite of even higher-resolution simulations, we decided to impose more conservative approach and apply a minimum cut-off in the number of particles that we would accept haloes to be sampled by. We did not impose a single cut-off, but tailored these according to our aims. To be more specific, we adopt the a particle cut-off of more than 700 particles for the case when we study the absolute value for the mass function and its evolution with redshift, which restricts the systematic bias to below $2\%$ according to Eq. (17). By contrast, when looking at the relative comparison between our variations in the cosmological model and the fiducial model, we are more tolerant and impose a cut-off at 150 particles (below $4\%$ according to Eq. (17). This is due to the fact that, this systematic bias should be, roughly speaking, cancelled or at the very least significantly reduced, when we inspect the ratios of $\nu f(\nu)$ measured from different models. In addition, we also choose to discard the highest mass bins, removing those with fewer than 100 haloes per bin. This limits the Poisson errors to better than $10\%$.
Figure 1: Top panel: Comparison between the measured ${\rm d}n/{\rm d}\log{M}$ as a function of halo mass. The data points show the results from the average of the 10 fiducial cosmology simulations, with the triangles, squares, circles, stars, and diamonds denote the results from $z=0.00$, $0.51$, $0.99$, $1.50$, and $2.05$. The lines denote the empirical models, where the dotted line denotes Sheth et al. (2001), the dashed lines Courtin et al. (2011), and dash-dotted lines Watson et al. (2013). Bottom panel: Fractional difference of mass functions with respect to the model of Courtin et al. (2011). Although only one base line is shown here, the differences are actually with respect to the model at each of the redshifts. Points are as in top panel.
We next estimate the mass function ${\rm d}n/{\rm d}\log m$ using the expression
$\frac{\widehat{{\rm d}n}}{{\rm d}\log{m}}(M_{i})=\frac{M_{i}\times N_{i}}{V\times\Delta M_{i}}\ ,$ (18)
where $V$ is the volume of the simulation box, $\Delta M_{i}$ and $N_{i}$ are the bin width, and number counts of the haloes of the $i^{\rm th}$ bin, respectively. Courtin et al. (2011) found that the binned mass function is more sensitive to the bin width when using the averaged mass in each bin to represent $M_{i}$ than using the bin-centre mass. Thus here we represent $M_{i}$ with the bin-centre mass.
Figure 1 shows the measured mass functions ${\rm d}n/{\rm d}\log{m}$ as a function of halo mass and for several redshifts out to $z=2$. For comparison we also present results for the fitting functions of Sheth et al. (2001) (dotted line), Courtin et al. (2011) (dashed line), and Watson et al. (2013) (dash-dotted line). From the plots it is clear that the abundances evolve strongly with mass and redshift.
In Appendix A we investigate the effect of bin width on our binned mass functions, and find that binning leads to an overestimate of the mass function, especially towards high redshift and for the higher masses or equivalently rare peaks. As a result, we also develop a binning correction method, the details of which can be found there. However, in this work, since we primarily make use of relative differences in mass functions between different cosmologies rather than focus on the absolute value, we find that the binning correction factors also cancel in the ratios. Finally, we note that, if one wishes to accurately measure the absolute value of the halo mass function to high precision, then binning corrections should be taken into account when less than 10 bins per decade of halo mass are used.
Next, in order to estimate the multiplicity function $\nu f(\nu)$ we need to complete the right-hand-side of Eq. (8), i.e., determining the quantity ${\rm d}\log{m}/{\rm d}\log\nu$. To do this, we note that this may also be expressed in terms of $R$ and $\sigma^{2}$ in the following way:
$\frac{{\rm d}\log m}{{\rm d}\log\nu}=-3\frac{{\rm d}\log R}{{\rm d}\log\sigma^{2}},$ (19)
which depends only on the present linear matter power spectrum $P_{\rm m}(k)$. This can be conveniently estimated from Eq. (15) by either computing the numerical derivative or evaluating the expression:
$\frac{{\rm d}\log\sigma^{2}(R)}{{\rm d}\log R}=\int\frac{{\rm d}k^{3}}{(2\pi)^{3}}P_{\rm m}(k)W(kR)V(kR)\ .$ (20)
where $V(y)$ is defined as:
$V(y)\equiv 2\left[9y\cos y+3(y^{2}-3)\sin y\right]/y^{3}\ .$ (21)
Note that here we employ the same constraints on the limits of the integral as for the case of evaluating $\sigma(m)$.
Figure 2: Top panel: Comparison between the measured $\nu f(\nu)$ in the fiducial cosmology as a function of $\nu$. The points and lines are as the same as presented in Fig. 1. Bottom panel: Ratio of the data with respect to the universal form presented in Courtin et al. (2011). Points and lines are as before.
Figure 2 shows a simple comparison between the measured $\nu f(\nu)$ and the fitting functions in Sheth et al. (2001) (dotted line), Courtin et al. (2011) (dashed line), and Watson et al. (2013) (dash-dotted line). The $\nu f(\nu)$ at all 5 redshifts, i.e, $z=0.00$, $0.51$, $0.99$, $1.50$, and $2.05$ are measured from a total of 10 runs of the fiducial cosmology, amounted to a volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. It can be seen that for the regime $\nu\geq 5$ that all of the data points form a similar locus. However, for the regime $\nu<5$ the data for $z=0$ are roughly 5% higher amplitude than the results for higher redshift. However for these data they exhibit a redshift-independent universality. Thus for our fiducial simulations we see that universality can only be said to be approximately true. These findings are in agreement with earlier results from Crocce et al. (2010) and Courtin et al. (2011).
We also note that the Courtin et al. (2011) function describes our $z=0$ data to an accuracy of the order $\sim 5\%$, but is systematically too high for our higher redshift data. On the other hand, the Watson et al. (2013) function underestimates our low mass haloes at $z=0$ by $\sim 5\%$, but describes the small to medium-$\nu$ regime to high accuracy, though deviates strongly for the high $\nu$ case.
3.4 Estimate of the error on the HMF
We next estimate the errors on the HMF. As was first shown by Smith & Marian (2011) the errors on the binned estimates of the HMF measured in CDM simulations are not independent, but are in fact correlated. This correlation arises due to the fact that halo abundances are modulated by the large scale modes of the power spectrum. Owing to the fact that we only have 10 realisations of our fiducial model, and single runs for each of the variational models, to estimate this covariance, we exploit the estimator described in Section 5.3 in Smith & Marian (2011) (see also Crocce et al., 2010). The approach is to take advantage of the fair sample hypothesis and divide the total volume of all the simulations into a number of smaller cubes, which provides us with a larger number of quasi-independent volumes but each with a smaller subvolume: the number of quasi-independent realisations is simply given by $N_{\rm tot}=N_{\rm runs}\times N_{\rm sub}$ where $N_{\rm runs}$ is the number of simulation runs and $N_{\rm sub}$ is the number of subcubes of each run. Although these realisations are not perfectly independent, Smith & Marian (2011) checked that there is no significant impact on the measurement of the HMF covariance.
The covariance of the binned mass function estimates can then be estimated using:
$\mathbf{M}x_{ij}$ $\equiv\left\langle n_{i}n_{j}\right\rangle-\left\langle n_{i}\right\rangle\left\langle n_{j}\right\rangle$
$=\bigg{(}\frac{N_{\rm sub}}{V_{\rm tot}}\bigg{)}^{2}\frac{1}{\Delta M_{i}\Delta M_{j}}\frac{1}{N_{\rm tot}}\sum_{\alpha,\beta=1}^{N_{\rm tot}}N_{\alpha,i}N_{\beta,j}-n_{i}n_{j},$ (22)
where $n_{i}=N_{i}/(V_{\rm tot}\times\Delta M_{i})$ is the estimate of the mass function in the $i$th bin averaged over all subcubes over all realisations, $V_{\rm tot}$ is the total volume of all runs, $N_{\alpha,i}$ and $N_{\beta,j}$ are the counts of the $\alpha^{\rm th}$ subcube in the $i^{\rm th}$ bin and the counts of the $\beta^{\rm th}$ subcube in the $j^{\rm th}$ bin, respectively. The covariance of the $\nu f(\nu)$ estimates can then be obtained through repeated application of the factors from Eq. (8):
$\mathbf{N}_{ij}$ $=\frac{m_{i}^{2}m_{j}^{2}}{\overline{\rho}^{2}_{0}}\mathbf{M}_{ij}\left|\frac{{\rm d}\log{m}}{{\rm d}\log{\nu}}\right|_{m_{i}}\left|\frac{{\rm d}\log{m}}{{\rm d}\log{\nu}}\right|_{m_{j}}\ .$ (23)
Finally, to examine the strength of the bin-to-bin covariance it is useful to compute the correlation matrix:
$\mathbf{r}_{ij}=\frac{\mathbf{M}_{ij}}{\sqrt{\mathbf{M}_{ii}\mathbf{M}_{jj}}}\ ,$ (24)
whose diagonal elements are unity and whose off diagonal terms are bounded to $\mathbf{r}_{ij}\in\left[-1,1\right]$. Note that the resultant correlation matrix is identical if we use either $\mathbf{M}_{ij}$ or $\mathbf{N}_{ij}$.
Figure 3: Fractional errors on ${\rm d}n/{\rm d}\log{M}$ at $z=0$ for the fiducial model. The results for $N_{\rm runs}=1$ (using only the fiducial run) and $N_{\rm runs}=10$ (using 10 runs of the fiducial model) are shown here. The blue, orange, green and red circles denote the errors estimated using the number of divisions of 3, 4, 5 and 6, respectively, on each side of simulation box, for $N_{\rm runs}=1$. The purple, brown, pink and grey triangles denote those for $N_{\rm runs}=10$. The 2 black dashed lines denote the fractional Poisson error for both cases. The 2 dotted horizontal lines mark the fractional errors of $10\%$ and $1\%$. The errors weakly depend on $N_{\rm sub}$ regarding the same $N_{\rm runs}$.
Figure 3 presents the fractional errors on the halo mass function estimator, i.e. $\sqrt{\mathbf{M}_{ii}}/n_{i}$. Here we show results obtained for the errors for the fiducial model at $z=0$ when we have computed the estimate from Eq. (22) with $N_{\rm sub}=\{3^{3},4^{3},5^{3},6^{3}\}=\{27,64,125,216\}$. In the plot the circled points present the estimates obtained for the first run of fiducial simulations and triangles give the results for the combination of all of the 10 fiducial runs. We see that for a single run the fractional errors are $\lesssim 1\%$ for the mass range $M\lesssim 3\times 10^{12}\ h^{-1}{\rm M}_{\odot}$, but are $\lesssim 10\%$ over the whole mass range considered. On the other hand, when averaging over all 10 realisations our error estimates are about 3 times smaller and are $\lesssim 1\%$ for haloes with $M\lesssim 1\times 10^{14}\ h^{-1}{\rm M}_{\odot}$.
The figure clearly demonstrates that the errors depend only weakly on our choice for $N_{\rm sub}$ over the whole mass range. In what follows we choose $N_{\rm sub}=64$ to compute the errors for later analysis, which corresponds to a size of sample box of $L_{\rm box}=125\ h^{-1}\rm{Mpc}$. In addition, the estimated error agrees well with the fractional Poisson error (dashed lines) at the high-mass end where the sample size is relatively small. As the number of haloes grows towards the low-mass bins, the error becomes dominated by sample variance, which leads to an increase of statistical error over the Poisson error (Crocce et al., 2010; Smith & Marian, 2011).
Figure 4: Correlation matrix of the HMF estimated using the sub-cube algorithm described in §3.4 with $N_{\rm sub}=64$ for a total of 10 runs in the fiducial cosmology at $z=0$.
Figure 4 shows the correlation matrix $\mathbf{r}_{ij}$ estimated using all 10 runs in the fiducial cosmology at $z=0$, which amounts to a total volume of $1.25\ h^{-3}\rm{Gpc}^{3}$. The estimate is constructed using $N_{\rm sub}=64$ per simulation. The figure clearly shows that the estimates of low- to intermediate group-mass bins are highly covariant, i.e., the cross-correlation coefficient $\mathbf{r}_{ij}\gtrsim 0.5$ for haloes with $M\lesssim\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). The correlation matrix only becomes close to diagonal towards the highest mass halo bins with $M\sim 10^{15}\ h^{-1}{\rm M}_{\odot}$. Each of the 10 runs has a similar correlation matrix but noiser, the results of which can be found in the Appendix B.
Figure 5: Comparison of ${\rm d}n/{\rm d}\log{M}$ as a function of halo masses at $z=0$. The top panels show the ${\rm d}n/{\rm d}\log{M}$ measured from the variational runs (coloured solid lines) and fiducial run (black solid line), and the other panels show the relative difference between them. The coloured squares and triangles denote the data for the positive and negative models respectively, regarding each cosmological parameter. The grey area marks the relative difference of ±$\pm 5\%$.
Figure 6: Comparison of $\nu f(\nu)$ as a function of $\nu$ at $z=0$. The top panels show the $\nu f(\nu)$ measured from the fiducial run (black points) and variational runs (coloured points), as well as the assumed universal locus (black dash-dotted line) estimated using linear interpolation. The two black dotted vertical lines delineate the most left and right black points, the locus beyond which is considered to be invalid. The rest panels show the relative difference between variational runs and the locus. The coloured squares and triangles represent the data for respectively the positive and negative variations of each of the cosmological parameters. The grey area marks the relative difference of ±$\pm 5\%$. For all variational models, most data points reside within the grey area except a few at $\nu\gtrsim 5$, which shows an approximate universality of $\nu f(\nu)$ at $z=0$ to a different extent exists with respect to the considered cosmological parameters.
3.5 Testing the universality of $\nu f(\nu)$
As was shown in §3.3 universality only approximately holds for our simulations. However, one of the further aims of this work is to explore to what extent $\nu f(\nu)$ is approximately universal and the range of cosmologies for which it can hold, especially when confronted with extensions beyond $\Lambda$CDM.
To do this then we first measure $\nu f(\nu)$ for our fiducial suite of simulations at each epoch that we have selected as described in the previous sections. While averaging over all the simulations would give us the best estimate of the locus for each redshift, we do not do this. Instead we take only the results for our run 1 simulation, since its random phase field matches that of our variation simulations and so features that can be attributed to large-scale cosmic variance can, to some extent, be mitigated in the relative difference. At each epoch, we therefore set up a linear interpolator function of the measured fiducial locus for $\nu f(\nu)$ in log-log scale. We then measure the $\nu f(\nu)$ in the variational runs and compare these data with the corresponding value obtained from the interpolated locus. In this way we are able to determine to what extent the cosmological variations can be said to hold to the locus from the fiducial run, at that epoch, respectively.
When comparing the data for the variations with the fiducial locus we apply a reduced chi-square test statistic to quantify the extent of overall universality for each of the variational runs at different redshifts. This is given by:
$\chi_{n}^{2}=\frac{\chi^{2}}{n}\quad\mathrm{where}\quad\chi^{2}=\sum_{i=1}^{N}\bigg{[}\frac{\Delta\nu f(\nu)}{\sigma}\bigg{]}_{i}^{2}\ ,$ (25)
where $\Delta\nu f(\nu)$ is the difference between the value of $\nu f(\nu)$ measured in the variational run and the interpolated locus from the fiducial run. $\sigma$ is the estimated error, $n$ is the number of degrees of freedom, i.e., the number of data points minus the number of fitted parameters. In our case, we do not include any parameters in the function of the locus. The reduced chi-square is usually used to describe the goodness-of-fit, with a value of $\chi_{n}^{2}\lesssim 1$ indicating a good fit since the deviations of the data are about the same or smaller than the size of their errors.
4 Results
In this section, for reference, we present the results for ${\rm d}n/{\rm d}\log{M}$ at $z=0$ for the variational and fiducial runs. We then present the results of $\nu f(\nu)$ measured from these runs at the same epoch. Finally, we include the results of $\nu f(\nu)$ at higher redshifts to show how the universality evolves with redshift honing in on the response with changes to each of the cosmological parameters varied.
4.1 Comparison of ${\rm d}n/{\rm d}\log{M}$ at $z=0$
Figure 5, top panels, shows the absolute mass function ${\rm d}n/{\rm d}\log{M}$ estimated for the variational and fiducial runs at $z=0$. The left panels show the impact for variations in the dark energy density parameter $\Omega_{\rm DE}$ and the equation of state parameters $w_{0}$ and $w_{a}$, along with the variations due to the physical density of CDM $\omega_{\rm c}$. The right panel shows the variations with the power spectral amplitude $A_{\rm s}$, its spectral index $n_{\rm s}$ and its running $\alpha_{\rm s}$, along with the physical density of baryons $\omega_{\rm b}$. The bottom panels show the relative difference between each of the variations and the fiducial model for a given cosmological parameter. We find that, at this redshift, the largest differences are for the variation of $\Omega_{\rm DE}$ (equivalently $\Omega_{\rm m}$), which are greater than $15\%$ and span nearly the whole mass range ($2\times 10^{11}\leq M\leq 5\times 10^{14}\ h^{-1}{\rm M}_{\odot}$). For other models, the deviations are almost within $5\%$ (grey area) for $M\lesssim 8\times 10^{13}\ h^{-1}{\rm M}_{\odot}$, but beyond this mass range growing differences can be observed.
As expected, the variations of some parameters affect ${\rm d}n/{\rm d}\log{M}$ in a very similar way but with differing strength. For example, the variations with respect to $w_{0}$, $w_{a}$, $\omega_{\rm c}$, $\omega_{\rm b}$, and $A_{\rm s}$ all lead to slightly positive/negative deviations for small mass scales, and this trend gradually reverses when considering larger mass scales. By contrast, the variations with respect to $n_{\rm s}$ and $\alpha_{\rm s}$ appear to affect the mass functions on all mass-scales at a similar level. Specifically, we find the positive/negative variation of them entirely increases/decreases the halo abundance over the whole mass range. Overall, it can be seen that the mass function, when expressed in the form of ${\rm d}n/{\rm d}\log{M}$, is quite sensitive with respect to these parameters, and it is difficult to accurately quantify the effect of the varied parameters on ${\rm d}n/{\rm d}\log{M}$. This statement is strengthened when one considers the evolution of the halo abundances with redshift.
Figure 7: The same as relative difference in Fig. 6 but for more selected redshifts between $z=0$ and $6.75$. Each of the columns is for the positive (squares) and negative (triangles) variations of the same cosmological parameter, with an increase of redshift from the top down. The corresponding redshifts are shown on the right side of each row. It can be seen that an increase of redshift does not impact the approximate universality observed at $z=0$ (shown in Fig. 6). The $\nu f(\nu)$ plot for these redshifts can be found in the Appendix C (Fig. 11).
4.2 Comparison of $\nu f(\nu)$ at $z=0$
Figure 6 presents the $z=0$ mass functions recast in the form of $\nu f(\nu)$, with estimates obtained in accordance with the discussion in §3.3. The top panels show measurements from the fiducial (black circles) along with the linear-interpolation model fit to this data (dash-dotted lines). The results for the variations in cosmology are presented as the coloured circles. The other panels show the relative difference between the variational runs and the linear-interpolation model of the fiducial data. The figure clearly shows that the relative differences between the variational results and the fiducial model are almost all within ±$\pm 5\%$ over the whole range of $\nu$ except a few points in the large-$\nu$ tail. However, here the errors are dominated by the low-counts in the bins, with Poisson errors being $\Delta\nu f(\nu)/\nu f(\nu)\lesssim 10\%$, given our cut-off of 100 haloes per bin.
It can also be seen in the $\nu\lesssim 0.5$ regime that the ratios for all variational models are slightly shifted from the universal locus by $\lesssim 2\%$. A possible explanation for this follows from the discussion in §3.3: the lower-mass haloes are sampled by a smaller number of particles and consequently their masses are overestimated (Warren et al., 2006). In order to investigate this in more detail, we ran additional 4 simulations with different resolutions but maintaining all other code parameter settings the same, i.e., same box size $L_{\rm box}=500\,h^{-1}{\rm Mpc}$ but different particle numbers $N=2048^{3},1024^{3},512^{3},$ and $256^{3}$. We found that the bias on the small-$\nu$ tail of $\nu f(\nu)$ can be significantly removed by using successively higher-resolution simulations. This implies the possibility that the universality holds to greater precision for these low-$\nu$ values than shown in the figure and that it could also possibly be extended to even smaller-mass haloes. We reserve that investigation for future work.
We also note that there appears to be features in the shape of the $\nu f(\nu)$ values estimated from all of the variational simulations and these features become more accentuated as one considers the $\nu>1$ data. We have checked that the appearance of these features can be ascribed entirely to the finite volume effects and the specific phase distribution of the initial Gaussian random field in Fourier space. On averaging over all 10 of the fiducial runs the $\nu f(\nu)$ locus that we recover is much smoother and for each realisation displays different features. However, when we compare the variation models to the fiducial model, it was important not to use this averaged $\nu f(\nu)$ spectrum since in this way the sample variance effects can be better mitigated for. In Appendix C we also show that the features are suppressed when one considers the $\nu f(\nu)$$\nu$ locus as a function of redshift (see Fig. 11).
In Table 2 we present a more quantitative analysis of the data and compute the reduced-$\chi_{n}^{2}$ of the variational data as described in Eq. (25). As the table shows, we find that for all cosmological variations considered at $z=0$ the reduced-$\chi_{n}^{2}$ is $\lesssim 1$ (the largest deviation being for the negative variation in $A_{\rm s}$). This indicates that the mass function can be said to possess a universal form over the mass range $(2\times 10^{11}h^{-1}{\rm M}_{\odot}<M<7\times 10^{14}h^{-1}{\rm M}_{\odot})$. We also found that the values of $\chi_{n}^{2}$ could be reduced by up to a half of the values quoted if we tightened the low-mass cutoff mass to include only haloes with 700 particles or more.
Table 2: Reduced Chi-square $\chi_{n}^{2}$ for each of the variational runs at different redshifts. Except the first column that denotes redshift, every two columns denote the positive and negative variations of each cosmological parameter. A value of $\chi_{n}^{2}\lesssim 1$ indicates a good universality of $\nu f(\nu)$.
$z$ $w_{0}+\Delta$ $w_{0}-\Delta$ $w_{a}+\Delta$ $w_{a}-\Delta$ $\Omega_{\rm DE}+\Delta$ $\Omega_{\rm DE}-\Delta$ $\omega_{\rm c}+\Delta$ $\omega_{\rm c}-\Delta$
$0.00$ $0.966$ $0.949$ $0.483$ $0.240$ $0.576$ $0.785$ $0.592$ $0.862$
$0.51$ $1.949$ $0.783$ $0.633$ $0.611$ $0.313$ $0.403$ $0.933$ $1.166$
$0.99$ $0.260$ $1.448$ $1.149$ $0.924$ $0.567$ $0.856$ $0.513$ $1.137$
$1.50$ $0.142$ $0.745$ $0.436$ $1.235$ $0.737$ $1.420$ $0.614$ $1.627$
$2.05$ $0.170$ $0.221$ $0.232$ $0.133$ $0.852$ $0.947$ $2.246$ $1.059$
$3.08$ $0.105$ $0.114$ $0.115$ $0.048$ $0.708$ $0.874$ $1.997$ $0.993$
$4.54$ $0.057$ $0.066$ $0.067$ $0.125$ $0.269$ $0.439$ $0.753$ $0.548$
$6.75$ $0.118$ $0.291$ $0.062$ $0.178$ $0.233$ $0.220$ $0.641$ $0.246$
$z$ $\omega_{\rm b}+\Delta$ $\omega_{\rm b}-\Delta$ $A_{\rm s}+\Delta$ $A_{\rm s}-\Delta$ $n_{\rm s}+\Delta$ $n_{\rm s}-\Delta$ $\alpha_{\rm s}+\Delta$ $\alpha_{\rm s}-\Delta$
$0.00$ $0.107$ $0.139$ $0.517$ $1.122$ $0.422$ $0.676$ $0.097$ $0.283$
$0.51$ $0.190$ $0.328$ $0.740$ $1.278$ $0.596$ $0.899$ $0.195$ $0.407$
$0.99$ $0.417$ $0.308$ $1.304$ $1.569$ $1.080$ $0.824$ $0.435$ $0.333$
$1.50$ $0.393$ $0.191$ $2.205$ $1.605$ $0.800$ $0.759$ $0.341$ $0.187$
$2.05$ $0.110$ $0.139$ $0.491$ $1.313$ $0.441$ $0.563$ $0.213$ $0.148$
$3.08$ $0.118$ $0.153$ $0.620$ $1.007$ $0.431$ $0.546$ $0.199$ $0.123$
$4.54$ $0.129$ $0.129$ $0.293$ $0.415$ $0.525$ $0.309$ $0.220$ $0.264$
$6.75$ $0.101$ $0.215$ $0.470$ $0.375$ $0.498$ $0.498$ $0.265$ $0.422$
4.3 Comparison of $\nu f(\nu)$ at higher redshifts
In Figure 7 we explore the redshift evolution of the universal multiplicity function $\nu f(\nu)$ ratio of the variational runs with respect to the fiducial model. The top panel shows results for the parameters $w_{0}$, $w_{a}$, $\Omega_{\rm DE}$ and $\omega_{\rm c}$, whereas the bottom panel shows results for $\omega_{\rm b}$, $n_{\rm s}$, $A_{\rm s}$ and $\alpha_{\rm s}$. Going from top to bottom the subpanels show the $\nu f(\nu)$ ratio results for the redshifts $z=\{0.0,0.51,0.99,1.5,2.05,3.08,4.54,6.75\}$. We notice that the approximate universality observed at $z=0$ holds to better than $\lesssim 5\%$ for higher redshifts. The possible outliers being for $A_{\rm s}$ and $\omega_{\rm c}$. These qualitative observations are more quantitatively demonstrated in the results from Table 2. This shows that most models share a similar tendency that as redshift goes up, the values of $\chi_{n}^{2}$ firstly increase and then start to drop at $z>1$ or $1.5$, which possibly indicates that universality holds better at low and high redshifts.
In more detail, for the parameters $\omega_{\rm b}$ and $\alpha_{\rm s}$ we find that the $\nu f(\nu)$ locus at higher redshifts is incredibly well described by the locus from the fiducial model. Looking at the simulations that vary the dark energy-related parameters $w_{0}$ and $w_{a}$, we again see good agreement for all relevant redshifts but with a potential deviation for $0.5\lesssim z\lesssim 1.5$, where the universality appears to be slightly broken. At higher redshifts this deviation is no longer visible. For $\Omega_{\rm DE}$, or equivalently $\Omega_{\rm m}$, the points become very close to the locus in the small- and medium-$\nu$ regimes at $z\gtrsim 0.5$, whereas moderate deviations can be observed with the large-$\nu$ tail at each of the redshifts. Likewise, regarding $\omega_{\rm c}$ and $A_{\rm s}$, a relatively good match can be seen with the small and medium scales of $\nu$ at all redshifts, but the points in the large-$\nu$ regime severely deviate from the locus. Interestingly, a similar trend of deviations can still be observed at some of the higher redshifts for both parameters. It is very likely that these variations have an equivalent influence on $\nu f(\nu)$. looking at $n_{\rm s}$, we see a good match to the fiducial locus, but again with some possible deviations around $z=1$.
5 Discussion
5.1 Utility of a universal FoF halo mass function
In a number of studies focused on cluster cosmology the utility of the FoF HMFs has been foregone in favour of studying haloes identified through either the spherical overdensity algorithm (Tinker et al., 2008; McClintock et al., 2019; Costanzi et al., 2019a, b). The objections mostly focus on the notion that the haloes selected by a spherical overdensity algorithm are more physically meaningful quantities since one can link the objects found to the virialisation overdensities resulting from the spherical collapse model, whereas the FoF algorithm with fixed $b=0.2$, links to haloes of varying overdensity (Courtin et al., 2011; More et al., 2011). Further, it is speculated that the cluster observable–halo mass distribution function can be more readily calibrated. These arguments can be countered by noting that it is well known that the condensations found in simulations are not spherical (for example see Courtin et al., 2011; More et al., 2011). In fact gravitational collapse leads to a distribution of prolate and oblate ellipsoids (Bardeen et al., 1986; Sheth et al., 2001; Jing & Suto, 2002). In this sense, one should consider spherical haloes as being less physically meaningful. Treating ellipsoidal haloes as spherical will lead to underestimates of the mass. Secondly, owing to the fact that the cluster observable–halo mass relation requires some detailed knowledge of galaxy formation physics and observational selection effects, there is no one-to-one correspondence between any 3D halo identified in simulations and any cluster observable. These have to be fully calibrated. With the FoF algorithm, one can at least be sure that one is not missing mass, which can happen in SO identified haloes.
Ultimately, for accurate cluster cosmology, one simply needs to carefully model either of the conditional probability density functions $P(X|M_{\rm SO})$ or $P(X|M_{\rm FoF})$, where $X$ is a given cluster observable. However, when it comes to making accurate predictions for the cluster counts, these will be obtained by convolving the above PDFs with the respective model for the mass functions. Our findings in this work suggest that the modelling of the $mn(m)$ for FoF haloes is simplified by the ability to exploit the near universality shown, especially when one considers extensions beyond the simple $\Lambda$CDM model. All of this leads us to the recommendation that it is $P(X|M_{\rm FoF})$ one should expend resources to calibrate over other choices.
Figure 8: The values of $\sigma_{8}-\Omega_{\rm m}$ covered by all snapshots between $z=0$ and $6.75$ of all cosmological models used in this work. Red, blue, and green transparent circles respectively denote the positively, negatively varied models, and the fiducial model. The points of each model starts from upper left with $z=0$ to lower right with $z=6.75$. The grey area shows the best-fit $\sigma_{8}-\Omega_{\rm m}$ degeneracy obtained from cluster abundance in Abdullah et al. (2020). A good overlap can be seen with $\Omega_{\rm m}\lesssim 0.4$, which indicates where our results could apply to in the $\sigma_{8}-\Omega_{\rm m}$ plane.
5.2 Parameter space coverage
It is well known that when constraining the cosmological parameters from the HMF there is a degeneracy between $\sigma_{8}$$\Omega_{\rm m}$. It is therefore interesting to examine where on this plane our results of near universality hold. To do this, we can think of the results from higher redshift snapshots as defining a different world model with a higher $\Omega_{\rm m}$ content and a corresponding lower $\sigma_{8}$. Thus we can use these results to cover a wider range of locations on the $\sigma_{8}$$\Omega_{\rm m}$ plane than one would expect from simply examining the $z=0$ models alone.
Figure 8 shows the collection of points in the $\sigma_{8}$$\Omega_{\rm m}$ plane that are covered by 47 of the snapshots that we analysed between $z=0$ and $6.75$ in all of the cosmological models used in this work. In the plot we also compare our parameter space coverage with the best-fit $\sigma_{8}-\Omega_{\rm m}$ degeneracy obtained from cluster abundance in Abdullah et al. (2020). We see that there is a good overlap between our sampled points and this curve for $\Omega_{\rm m}\lesssim 0.4$. However, we expect that our results should hold well beyond the locus of points shown. It is our intention to explore this further in a future study.
5.3 Caveats to the analysis
Some caveats to over-interpreting our findings are that our approach of examining the ratios has neglected to account for the noise in the fiducial model and the errors that are shown are those for the variational models alone. Furthermore, in our analysis of the reduced-$\chi^{2}$ as described in Eq. (25), we are also neglecting the errors in the interpolated locus of the fiducial model. A better approach would be to take account of these errors when performing the test. This might possibly be done by building a Gaussian process model for the fiducial model and then use that to forward model predictions and errors for the actual bins used in the variational models (see McClintock et al., 2019, for discussion of the emulation). A further source of improvement would be to consider the inclusion of the off diagonal components of the covariance matrix. However, this would likely require a larger suite of simulations in order to minimise the errors induced in computing the inverse of the matrix. We leave these investigations for future work.
6 Conclusions
In this paper we have investigated the universality of the abundance of dark matter haloes in the $\Lambda$CDM and $w$CDM frameworks and have also explored the impact of modifications of the primordial physics through variations with respect to the amplitude, spectral index of the primordial power spectrum, and along with a possible running of the primordial power spectral index. For our exploration, we made use of the Dämmerung simulations, a suite of high-resolution ($N=2048^{3}$) dark matter only cosmological $N$-body simulations, carried out by Smith & Angulo (2019). The simulations consisted of 10 runs of a fiducial model that adopted the Planck 2014 cosmology and 16 variational runs that positively and negatively varied each of the 8 cosmological parameters around the fiducial model. We selected haloes using the FoF($b=0.2$) algorithm. This choice owed to the fact that previous studies had demonstrated a near universality for these objects and due to the advantage of computational efficiency and scalability.
In order to test the universality $\nu f(\nu)$ we built a linear interpolation model of the results from our $z=0$ fiducial run. On comparison of the $\nu f(\nu)$ measured from the variational runs with this model, we were able to assess the validity of the putative universal behaviour with respect to the 16 variations. We also selected a number of snapshots with a coverage of redshift $z\in[0,6.75]$ to inspect evolution with redshift. To make quantitative statements, we performed a reduced-$\chi^{2}$ test, for which we applied the subcubing approach introduced in Smith & Marian (2011) to compute the covariance matrices of the HMFs and so estimate statistical errors. We summarise our findings as follows:
  • Considering the resolution and volume of each of the Dämmerung simulations, we found that reliable results were obtained for the haloes in the mass range $(1.0\times 10^{12}h^{-1}{\rm M}_{\odot}\leq M\leq 7.0\times 10^{14}h^{-1}{\rm M}_{\odot})$. Comparing to measurements of ${\rm d}n/{\rm d}\log{M}$ from our fiducial model, we found that at $z=0$ the variations of the cosmological parameters $w_{0}$, $w_{a}$, $\omega_{\rm c}$, $\omega_{\rm b}$ and $A_{\rm s}$ mainly impacted ${\rm d}n/{\rm d}\log{M}$ for $M\gtrsim 10^{14}\ h^{-1}M_{\odot}$, whereas for mass scales $M\lesssim 10^{14}\ h^{-1}M_{\odot}$ the effects were more modest. On the other hand, the variations of $\Omega_{\rm DE}$ ($\Omega_{\rm m}$) led to strong changes in ${\rm d}n/{\rm d}\log{M}$ over the whole mass range. By contrast, variations arising due to modifications of $n_{\rm s}$ and $\alpha_{\rm s}$, led to moderate effects on all scales.
  • For the redshift range $(0\lesssim z\lesssim 6.75)$, we found that the relative differences between the $\nu f(\nu)$ across all variations at each redshift were better than ±$\pm 5\%$ over the whole range of $\nu$ considered, except for a few points in the large-$\nu$ tail. However, these excursions were likely due to sample variance. We applied the reduced-$\chi_{n}^{2}$ test to quantify the extent of the overall universality and found for the $z=0$ data values of $\chi_{n}^{2}\lesssim 1$, indicating good approximate universality at that redshift across the models we considered. For higher redshifts, the reduced-$\chi_{n}^{2}$ slightly increased and then dropped for $z\gtrsim 1.5$, which suggested that approximate universality holds more strongly at both low and high redshifts. However, some of the deviations we noted could be due to the fact that we did not fold in the sample and Poisson errors of the fiducial model into the ratio, which softens our conclusions at these redshifts and leaves room for further improved agreement between models.
  • Amongst all of the parameter variations that we considered, the universality of $\nu f(\nu)$ was found to hold to very high precision for $\omega_{\rm b}$ and $\alpha_{\rm s}$ across all $\nu$ and $z$ considered. The variations of all of the other parameters $w_{0}$, $w_{a}$, $\Omega_{\rm DE}$, $\omega_{\rm c}$, $n_{\rm s}$ and $A_{\rm s}$ was similarly good at low and high redshift, but with hints of a breakdown occurring at intermediate redshifts. However, all departures were $\lesssim 5\%$.
  • We noticed that for nearly all of the cosmological models considered the points in the small-$\nu$ regime (corresponding to haloes with mass scales $M\lesssim 10^{12}h^{-1}{\rm M}_{\odot}$) were slightly shifted from the universal locus. We concluded that this was likely caused by the systematic bias in the mass estimates for FoF haloes sampled by finite numbers of particles. This implied that the good approximate universality found here could actually be extended to an even wider range of halo masses. However, to confirm this, would require a suite of $N$-body simulations of even higher resolution and so we leave confirmation of this to future work.
In recent years there has been a rapid increase in the methods of identifying and classifying dark matter haloes in numerical simulations and in observational data. A number of theoretical studies have followed the path of studying haloes identified through the SO algorithm. However, it has been widely confirmed that the HMFs for these objects are far from a universality and thus require a large amount of simulations to accurately calibrate their abundance as a function of cosmological model and over time. By contrast, haloes identified through the FoF algorithm exhibit approximate universality with respect to a wide set of cosmological parameters. The advantage of this is that only several mass functions need to be accurately parameterised so that they can be deployed for likelihood modelling in cluster abundance studies.
In this work, we have completed a first step to producing an improved model for the HMF. We have quantitatively shown that FoF haloes do have the potential for building such models with the minimum required sets of numerical simulations. Our next step will be searching for an effective form of the universal $\nu f(\nu)$ that can provide percent-level constraints. The ultimate goal, of course, is to produce a theory forward modelling pipeline for modelling the likelihood function for observing cluster counts binned in terms of observable mass proxies for upcoming galaxy surveys such as Euclid. To achieve this, one needs to understand the conditional probability density of obtaining the observable given the underlying halo.
In this study we have also neglected how modifications of the baryonic physics may impact the universality of the mass function. Nor have we included the study of massive neutrinos or modifications of the dark matter particle modelling. These we shall reserve for future study.
Acknowledgements
RES acknowledges support from the Science and Technology Facilities Council (grant ST/X001040/1). YL thanks Baojiu Li for useful discussion. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We acknowledge that the results of this research have been achieved using the PRACE Tier-0 Research Infrastructure resource SuperMuc based in Garching Germany at the Leibniz Supercomputing Centre (LRZ) under project number 2012071360.
Data Availability
Simulation data used in this work can be made available upon request to the authors.
Appendix A Effect of binning on multiplicity function
We quantified the systematic bias on the binned mass function regarding bin width. Firstly, we assume some forms of the fitting functions as true mass functions, which are obtained by a HMF calculator hmf (Murray et al., 2013). Then we adopt our fiducial cosmology to estimate the binned mass functions from the fitting functions by computing the numerical integral of each bin over its bin width. The relative differences between the binned $\nu f(\nu)$ and the fitting $\nu f(\nu)$ are shown in Fig. 9, which indicates that binning would cause an overestimate of $\nu f(\nu)$. This can be seen with all 3 selected models of fitting functions, i.e., Watson(FoF) (squares), Tinker08 (triangles), and Bhattacharya (circles), although the extent of the overestimate could be a little different regarding each model. The left, middle and right panels are for a choice of 5, 8, and 11 bins per decade, respectively, in logarithmic scale of halo masses. By comparing them, it is understood that the bias can be theoretically reduced by increasing the number of bins applied. However, in practice, small bin width could cause a problem of empty bins as well as very wiggled mass function when counting haloes in numerical simulations. Thus an optimal choice of bin width should be checked case by case. Additionally, it can be seen that the binned mass functions tend to be much more biased towards high redshift with the same choice of bin width. For example, an overestimate is below $2\%$ at $z=0$ with 8 bins per decade used, whereas it could reach a maximum of $6\%$ at $z=6.75$. Therefore, if one tries to measure an actual mass function with binning method, the correction of the bias respecting bin width should be seriously taken into account.
Figure 9: Relative difference of the binned $\nu f(\nu)$ to fitting $\nu f(\nu)$ as a function of $\nu$. The left, middle, and right panels are computed using, respectively, 5, 8, and 11 bins per decade in logarithmic scale of halo masses. The squares, triangles and circles denote the fitting functions of Watson(FoF), Tinker08, and Bhattacharya, respectively. The different colours of points correspond to the redshifts shown in the legend.
Appendix B Results of correlation matrices for individual runs of fiducial cosmology
Fig. 10 represents the results of correlation matrices of ${\rm d}n/{\rm d}\log{M}$ for each of the 10 runs of the fiducial cosmology at $z=0$.
Figure 10: Correlation matrices of ${\rm d}n/{\rm d}\log{M}$ for 10 individual runs of the fiducial cosmology at $z=0$.
Appendix C Results of multiplicity function for higher redshifts
Fig. 11 shows the results of $\nu f(\nu)$ at $0\leq z\leq 6.75$. The assumed universal loci are wiggled at low redshift and tend to be smooth towards high redshift. It can be seen that the $\nu f(\nu)$ measured from all models are very close to the universal locus at each of the redshifts, especially going to high redshift.
Figure 11: Comparison of $\nu f(\nu)$ as a function of $\nu$ at different redshifts. Each panel shows the $\nu f(\nu)$ measured from the fiducial run (black points) and variational runs (coloured squares or triangles), as well as the assumed universal locus (black dash-dotted line) estimated using linear interpolator. The two black dotted vertical lines delineate the most left and right black points, the locus beyond which is considered to be invalid. The corresponding redshifts are shown on the right side of each row. The measured $\nu f(\nu)$ from all models are very close to the loci, especially going to high redshift.